OpenStudy (anonymous):

1. Find x if ƒ(x) = 0. Use the graph of f below. Assume the entire function is graphed below.

OpenStudy (anonymous):

(a.) x = 0, x = 4 (b.) x = -2, x = 4 (c.) x = -2, x = 0 (d.) x = -3, x = 6

jimthompson5910 (jim_thompson5910):

where does the graph cross the x axis? the x axis is the horizontal axis

OpenStudy (anonymous):

cross as in intersect or pass?

jimthompson5910 (jim_thompson5910):

yes or touch

OpenStudy (anonymous):

-2, 0 and 4, 0

jimthompson5910 (jim_thompson5910):

put another way: what points are on the graph AND on the x axis at the same time

jimthompson5910 (jim_thompson5910):

(-2,0) and (4,0), yes

jimthompson5910 (jim_thompson5910):

you then list the x coordinates of each x-intercept

jimthompson5910 (jim_thompson5910):

and you're done

OpenStudy (anonymous):

so answer is C: -2, 0?

OpenStudy (anonymous):

please respond

jimthompson5910 (jim_thompson5910):

the points were (-2,0) and (4,0) the X COORDINATES of each point are: -2 and 4 so x = -2 or x = 4 will make f(x) = 0 true

jimthompson5910 (jim_thompson5910):

the answer is not -2, 0 because x = 0 doesn't make f(x) = 0 true when x = 0, f(x) = 6. Notice how the point (0,6) is on the graph

OpenStudy (anonymous):

so the answer is B: -2, 4?

jimthompson5910 (jim_thompson5910):

correct